OpenStudy (anonymous):

Please help with this calculus problem! f(x) = x / (x^2 - 9) I need to sketch this function without using technology and I need to find: i) Extremas ii) Points of inflection iii) Concavities

OpenStudy (anonymous):

Hi Turing :)

OpenStudy (turingtest):

1) extrema occur when f'(x)=0 as we discussed before

OpenStudy (turingtest):

Hi :)

OpenStudy (anonymous):

Ok, for the derivative, I got: \[-x ^{2}+9/(x ^{2}-9)^{2}\]

OpenStudy (anonymous):

I have no idea on how to solve for x in this one

OpenStudy (turingtest):

I don't think that's right I'm sorry I have to go.... work on that derivative: use the product rule and chain rule

OpenStudy (anonymous):

Ok, bye

OpenStudy (anonymous):

AccessDenied can you help?

OpenStudy (accessdenied):

yeah, im just going over the problem myself to make sure i have it straight

OpenStudy (anonymous):

Ok

OpenStudy (turingtest):

rewrite it as\[x(x^2-9)^{-1}\]and use the product rule and chain rule (quotient rule is for suckers!)

OpenStudy (anonymous):

LOL thats how I did it (quotient rule)

OpenStudy (turingtest):

product rule is easier in many cases, as you will see

OpenStudy (accessdenied):

i used quotient rule too, even though i normally do product rule with negative exp :(

OpenStudy (turingtest):

you can derive the quotient rule from the product rule in about 2 seconds

OpenStudy (turingtest):

\[(uv^{-1})'=u'v^{-1}-uv^{-2}v'=v^{-2}(u'v-v'u)=\frac{u'v-v'u}{v^2}\]therefor the Q rule is a waste of time in my opinion, but to each their own :)

OpenStudy (anonymous):

:'(

OpenStudy (turingtest):

I should write\[(\frac uv)'=(uv^{-1})'=u'v^{-1}-uv^{-2}v'=v^{-2}(u'v-v'u)=\frac{u'v-v'u}{v^2}\]

OpenStudy (anonymous):

Are you sure that my derivative is wrong?

OpenStudy (turingtest):

no I didn't check :P just looked wrong like I said gotta go!

OpenStudy (anonymous):

OK, Bye

OpenStudy (accessdenied):

its pretty close it might just be what you're writing to show it

OpenStudy (accessdenied):

like, do you mean for it to be this? \[ -\frac{x^2 + 9}{(x^2 - 9)^2} \]

OpenStudy (anonymous):

Thats exactly what I got

OpenStudy (anonymous):

does it make a difference if I apply the - sign to the x?

OpenStudy (accessdenied):

yeah, then its correct... you just had it written in a not-so-obvious way. :P

OpenStudy (anonymous):

so its -x^2?

OpenStudy (accessdenied):

the "-" at the outside would go through to both the x^2 and 9 \[ \frac{-x^2 - 9}{(x^2 - 9)^2} \]

OpenStudy (anonymous):

Oooohh

OpenStudy (anonymous):

Ok, now that we have the derivative, how do we solve for x?

OpenStudy (anonymous):

Can you also include steps :P

OpenStudy (accessdenied):

we have to set both the numerator and the denominator = 0 and try to get the x-value (if the numerator = 0, then the expression is 0; and if the denominator is 0, then the expression is undefined) -x^2 - 9 = 0 (x^2 - 9)^2 = 0

OpenStudy (anonymous):

I got 3 for the first one, I dont know how to do the second one

OpenStudy (accessdenied):

the first one, you wouldn't have any real solutions. -x^2 - 9 = 0 -x^2 = 9 x^2 = -9 <-- no real solution to that!

OpenStudy (anonymous):

Oh lol

OpenStudy (accessdenied):

the second one actually does have solutions. (x^2 - 9)^2 = 0 sqrt( (x^2 - 9)^2 ) = sqrt( 0 ) |x^2 - 9| = 0 absolute values not needed because |a| = 0 <=> a=0 x^2 - 9 = 0 x^2 = 9 sqrt( x^2 ) = sqrt( 9 ) |x| = 3 x = +- 3

OpenStudy (anonymous):

Ok I get that

OpenStudy (accessdenied):

except, to be extrema, the point has to be on the graph of the original function. x=3 and x=-3 are both undefined for the original function as well. So, we'll have no extrema for the graph.

OpenStudy (anonymous):

Ok

OpenStudy (accessdenied):

So, do you remember how to find the points of inflection?

OpenStudy (anonymous):

You need to find the concavities

OpenStudy (anonymous):

You find the second derivative and the inflection is at that point., right?

OpenStudy (accessdenied):

the points where the second derivative is 0 or undefined, and exists for the first derivative as well.

OpenStudy (anonymous):

So, there is no point of inflection?

OpenStudy (accessdenied):

have you tested the second derivative?

OpenStudy (anonymous):

No

OpenStudy (accessdenied):

itd be good to check the second derivative first before you assume it doesnt! :)

OpenStudy (anonymous):

I dont know how to take the derivative of this function

OpenStudy (accessdenied):

okay, i gotta go through it a second time myself. its not easy to do, but we can use the same technique as the first derivative. There's just more going on this time. :P

OpenStudy (anonymous):

Yeah, I already made so many mistakes :/

OpenStudy (accessdenied):

heres what wolfram says: http://www.wolframalpha.com/input/?i=derivative+of+-%28x%5E2+%2B+9%29%2F%28x%5E2+-+9%29%5E2 yeah, i keep making mistakes too. I need to try one more time!

OpenStudy (anonymous):

This is very confusing

OpenStudy (anonymous):

I cant get this second derivative...

OpenStudy (accessdenied):

okay, i finally got it after realizing that i could multiply things together for magic i'll link my work, if it possibly helps

OpenStudy (anonymous):

That is very difficult to read, but atleast you got the right answer :P

OpenStudy (accessdenied):

i used the product rule with negative exponents i changed it to exponent form on the first equals-sign second, i used product rule then, i brought the negative exponents back into standard i got common denominators on the first term with the second by multiplying both top and bottom by (x^2 - 9). I also multiplied the -4x through i then multiplied the 2x through so that i could brng the fractions together and add like-terms once i had it down in that form, i just factored out a 2x and removed the negative.

OpenStudy (anonymous):

OK=k

OpenStudy (anonymous):

Hold on, im still trying to read your work :P

OpenStudy (accessdenied):

i could probably rewrite it in latex, it'd just take another minute. lol

OpenStudy (anonymous):

On the fourth = sign, that line, whats the denominator for the first fraction?

OpenStudy (anonymous):

the exponent on that term

OpenStudy (accessdenied):

should be "(x^2 - 9)^3"

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

on the last line its -2x^3-5 and then what

OpenStudy (accessdenied):

-2x^3 - 54x

OpenStudy (anonymous):

ok I got it now

OpenStudy (anonymous):

So what do we do with this second derivative

OpenStudy (accessdenied):

we want to find the points where it is 0 or undefined, where the first derivative does exist. It has the same points where it is undefined as the first derivative, but we have a new 0. The numerator has "2x" as one term, which when x=0, makes the numerator 0.

OpenStudy (anonymous):

So this is undefined aswell right?

OpenStudy (accessdenied):

the second derivative is undefined for x=3 and x=-3, but since they don't exist on our first derivative either, its not a inflection point.

OpenStudy (anonymous):

So what should I write for the inflection point?

OpenStudy (accessdenied):

the inflection point is at x=0 since that is where the second derivative is 0.

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

So we have to equate the second derivative to 0 but when we do that we get the numerator at 0 right?

OpenStudy (anonymous):

Am I right?

OpenStudy (anonymous):

AccessDenied, are you there?

OpenStudy (turingtest):

yes, set the numerator equal to zero

OpenStudy (anonymous):

Alright now that I know that the inflection point is at zero, and I think we've determined that there are no concavities we just need to find the extremas

OpenStudy (accessdenied):

concavity is from testing the intervals around the inflection points. sorry, i had to reply to somebody else about stuff. D:

OpenStudy (anonymous):

So what did we determine from taking the first derivative?

OpenStudy (accessdenied):

the first derivative was to find if there were any extrema we check if the x-values that make the second derivative 0 or undefined in the first derivative to make sure that they are on the graph since otherwise, it wouldn't be an inflection point.

OpenStudy (anonymous):

And we determined that there were no concavities right?

OpenStudy (accessdenied):

hmm.. not sure how we should test for this one. seems like the concavity changes between the x=3 and x=-3 as well...

OpenStudy (accessdenied):

maybe TuringTest is more familiar with it. D:

OpenStudy (anonymous):

You said that +-3 was undefined for the original function

OpenStudy (accessdenied):

@TuringTest could you explain this? I'm definitely not remembering how this would be done. >.<

OpenStudy (anonymous):

First derivative set equal to zero = extreamas. Second derivative set equal to zero = points of inflection

OpenStudy (turingtest):

you just pick a point in the interval the concavity changes at \(x=\pm3\) so we get the intervals\[(-\infty,-3)(-3,3)(3,\infty)\]so check a point in each interval, and plug that point into \(f''(x)\) if \(f''<0\) it is concave down if \(f''>0\) it is concave up

OpenStudy (anonymous):

Isnt +-3 undefined for the original function though?

OpenStudy (turingtest):

so? that can still be a point where the concavity changes

OpenStudy (anonymous):

Oh

OpenStudy (anonymous):

OK

OpenStudy (turingtest):

they are asymptotes that often changes concavity from one side to the other

OpenStudy (anonymous):

Ok, so how should I write the answer for the concavities?

OpenStudy (accessdenied):

so, are they also considered "inflection points"?

OpenStudy (turingtest):

http://www.wolframalpha.com/input/?i=plot%20f(x)%20%3D%20x%20%2F%20(x%5E2%20-%209)&t=crmtb01 notice how the concavity changes between x=-3 and x=3 the inflection points...